OpenStudy (anonymous):

Is it possible to prove that a+sqrt(m)=b+sqrt(n) hence a=b and sqrt(m)=sqrt(n)

OpenStudy (anonymous):

\[a-b=\sqrt{n}-\sqrt{m}\]You are only considering the special case when both sides =0. Both sides could be 1, pi, 16, so in short, no.